OpenStudy (anonymous):

The mean of a list of 80 numbers is 230. If four numbers 145,156,210, and 255 are added, what is the mean of the list of numbers?

OpenStudy (anonymous):

\[Average(Mean) = \frac{Sum Of Numbers}{Total Numbers}\]

OpenStudy (anonymous):

So, you have given Mean = 230 Total Numbers are : 80 So, Sum Of Numbers = Average*Total Numbers = 230*80 = 18400 Now from this Sum, you know 4 numbers are:145, 156, 210 and 255 So, their sum is : 145 + 156 + 210 + 255 = 766 So, Sum Of List = Sum Of Numbers - 766 So, Sum of List = 18400 - 766 = 17634.. So, to find Mean: Mean = (Sum Of List)/(80 - 4) = 17634/76 = 232.02 Answer...